# Semiclassical measure of the propagation between two topological insulators

### v2

@@ Abstract
We study the propagation of a family of initial conditions in the presence of two topological insulators without magnetic field where the interface is a smooth connected curve not compact without boundaries. The solution is governed by an adiabatic modulation of a Dirac operator with a variable mass. We determine the evolution of the semiclassical measure of the solution with a two-scale Wigner measure method by reducing the Dirac operator to a normal form.
Keywords : propagation, Dirac operator, semiclassical analysis, Wigner measure, semiclassical measure valued in a separable Hilbert space, time dependent two-scale semiclassical measure.
@@
Acknowledgments. The author is very grateful to his advisors Clotilde Fermanian-Kammerer and Nicolas Raymond for many insightful discussions and valuable suggestions. The author acknowledges the support of the Région Pays de la Loire via the Connect Talent Project HiFrAn 2022 07750, and from the France 2030 program, Centre Henri Lebesgue ANR-11-LABX-0020-01.
@@ 1 Introduction
1.1 Topological insulators and Dirac type equations
Topological insulators are materials that are insulating in their bulk but support robust currents along their boundaries, meaning that electrons can only move along the surface of the material. The Dirac equation has been derived in the context of topological insulators for honeycomb structures [9, 8, 5] such as graphene [22, 18, 7]. From a global point of view, Dirac-type equations are often a simpler continuum macroscopic description of transport in a narrow energy band near the band crossing [2, 12, 24].
@@ 1.1 Topological insulators and Dirac type equations
In this paper we consider the following system of the Dirac evolution equation on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
@@
$\left\{\begin{array}[]{ll}\Big{(}\varepsilon\mathrm{D}_{t}+\mathrm{m}(x)\sigma_{3}+\varepsilon\mathrm{D}_{1}\sigma_{1}+\varepsilon\mathrm{D}_{2}\sigma_{2}\Big{)}\psi^{\varepsilon}(t,x)=0,&\forall(t,x)\in\mathbb{R}\times\mathbb{R}^{2},\\ \psi^{\varepsilon}(0,x)=\psi^{\varepsilon}_{0}(x),&\forall x\in\mathbb{R}^{2},\end{array}\right.$ (1)
@@
where ${\varepsilon}$ is a small positive parameter, called semiclassical parameter. The function $\psi^{\varepsilon}(t,\cdot)$ also denoted $\psi^{\varepsilon}_{t}$ is interpreted as the wave function of an electron at time $t$ in a material consisting in two topological insulators. The family ${\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}}$ is a uniformly bounded (in $\varepsilon$) family of ${\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ of initial wave functions of the electron. The function $x\mapsto\mathrm{m}(x)$ is a smooth real-valued function (i.e. $\mathrm{m}\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and all its derivatives are bounded. The matrices ${\left(\sigma_{j}\right)_{1\leqslant j\leqslant 3}}$ are the Pauli matrices
We shall use the notation
$\mathrm{H}_{\varepsilon}\coloneqq\begin{pmatrix}\mathrm{m}(x)&\varepsilon\mathrm{D}_{1}-\mathrm{i}\varepsilon\mathrm{D}_{2}\\ \varepsilon\mathrm{D}_{1}+\mathrm{i}\varepsilon\mathrm{D}_{2}&-\mathrm{m}(x)\end{pmatrix}.$
Then the equation (1) rewrites $\varepsilon\mathrm{D}_{t}\psi_{t}^{\varepsilon}+\mathrm{H}_{\varepsilon}\psi_{t}^{\varepsilon}=0$.
In this model, the function $\mathrm{m}$ splits the space into two parts, which are the two topological insulators. From a modeling point of view, the function $\mathrm{m}$ is positive on one of the topological insulators and negative on the other so the interface between the two topological insulators is defined by
$\mathrm{E}\coloneqq\left\{x\in\mathbb{R}^{2}\ |\ \mathrm{m}(x)=0\right\}.$
We assume the set $\mathrm{E}$ is a connected curve not compact without boundaries. From a physical point of view, the sign of $\mathrm{m}$ is the insulating phase of a material. This generates a unidirectional propagation along the interface $\mathrm{E}$. This propagation is one of the most important issue of physical applications of topological insulators and superconductors [2, 24], photonics [19, 22, 23], acoustics [21, 20] and fluid mechanics [13].
Our aim is to describe the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to equation (1) for initial conditions in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. At least, we can understand limits as $\varepsilon$ goes to 00 by constructing semiclassical defect measures, also called Wigner measures. Then, we can deduce various properties of this kind of measures and thereby indirectly recover information about asymptotics as $\varepsilon$ goes to 00.
@@ 1.2 Edge states generated by coherent states
The system (1) has already been studied in [1] where are performed precise computation. The edge mode computed is generated by a very specific shape of semiclassical wave packet.
In [6], it was proved that some terms of the solution to equation (1) are negligible in $\mathrm{L}^{\infty}$-norm as $\varepsilon$ goes to 00. It has been proven, for initial conditions, that semiclassical wave packets concentrate at a point on the $\mathrm{E}$ curve with no constraint on orientation. However, these dispersive terms are not necessarily negligible in the $\mathrm{L}^{2}$ norm. Let us recall these results.
Let $x_{0}\in\mathrm{E}$ and $t\mapsto x_{t}$ the solution to the ODE
$\left\{\begin{array}[]{l}\dot{x}_{t}=\frac{\nabla\mathrm{m}(x_{t})^{\perp}}{\left|\nabla\mathrm{m}(x_{t})\right|},\\ \left.x_{t}\right|_{t=0}=x_{0}.\end{array}\right.$
Let us define, for all $t\in\mathbb{R}$, $r_{t}\in\mathbb{R}_{>0}$ and $\theta_{t}\in\mathbb{R}$, such that $r_{t}\begin{pmatrix}-\sin\theta_{t}\\ \cos\theta_{t}\end{pmatrix}=\nabla\mathrm{m}(x_{t})$ and
$\vec{V}_{\theta_{t}}\coloneqq\begin{pmatrix}e^{\frac{-\mathrm{i}\theta_{t}}{2}}\\ -e^{\frac{\mathrm{i}\theta_{t}}{2}}\end{pmatrix}.$
Following [6], we define wave packet concentrated at $(x_{0},\xi_{0})\in\mathbb{R}^{4}$ as a $\varepsilon$-dependent function satisfying for all positive $\varepsilon$ and for all $x\in\mathbb{R}^{2}$,
@@
$\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}](x)\coloneqq\frac{e^{\frac{\mathrm{i}}{\varepsilon}\xi_{0}\cdot(x-x_{0})}}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
@@
with $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$. We say that $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$ is oriented along $\vec{V}\in\mathbb{C}^{2}$ if there exists $f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}\right)$ such that for all $x\in\mathbb{R}^{2},\vec{f}(x)=f(x)\vec{V}$. The parameter $\xi_{0}$ is interpreted as the impulsion (or the speed) of an electron of wave function $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$.
Theorem 1.1 ([1] Theorem 2 and [6] Theorem A.1).
Let $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$. If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
then there exists $T>0$, such that for all $t\in(0,T)$, uniformly for $\varepsilon\in(0,1]$,
$\psi_{t}^{\varepsilon}=\mathrm{WP}^{\varepsilon}_{x_{t},0}+\mathcal{O}_{\mathrm{L}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{-1/4}\right)+\mathcal{O}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{1/2}\right).$ (2)
where $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ is a wave packet concentrated at $(x_{t},(0,0))$, oriented along $\vec{V}_{\theta_{t}}$. Moreover, we have the following.
  1. (i)
    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}$, then the equality (2) holds for all $t>0$ without $\mathcal{O}_{\mathrm{L}^{\infty}}\left(\varepsilon^{-1/4}\right)$.
  2. (ii)
    If there exists $f\in\mathscr{S}\left(\mathbb{R},\mathbb{C}\right)$ such that for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
    $\psi_{0}^{\varepsilon}(x)\coloneqq\frac{r_{0}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{0}\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{0}},$
    where $\mathrm{R}_{\theta}$ is the rotation of angle $\theta$ in $\mathbb{R}^{2}$, then uniformly for $\varepsilon\in(0,1]$, for all $(t,x)\in\mathbb{R}\times\mathbb{R}^{2}$,
    $\psi_{t}^{\varepsilon}(x)=\frac{r_{t}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{t}\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{t}}\ +\ \mathcal{O}\left(\varepsilon^{1/2}\langle t\rangle\right).$
  3. (iii)
    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}^{\perp}$, then $\mathrm{WP}^{\varepsilon}_{x_{t},0}=0$.
The initial wave packet is concentrated on the curve $\mathcal{C}\coloneqq\mathrm{E}\times\{(0,0)\}\subset\mathbb{R}^{4}$ and so is the wave packet at time $t$. Physically, the wave packets under consideration have no impulse.
Let us informally describe these wave packets. In the point (ii), the variables $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}$ and $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}$ represent the tangent and the normal variable to the curve $\mathrm{E}$, respectively. They do not play the same role in the considered wave packet. In particular, for all $t\in[0,T)$, the wave packet $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ have a Gaussian behavior in the normal variable and is localized near the point $(x_{t},(0,0))$ in the phase space.
If the orientation is not along $\vec{V}_{\theta_{0}}$ or the decay with respect to the normal variable is different, then the $\mathrm{L}^{\infty}$-remaining part is not necessarily negligible in $\mathrm{L}^{2}$-norm. Therefore Theorem 1.1 does not give a full description of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
To prove this theorem, the Hamiltonian $\mathrm{H}_{\varepsilon}$ is reduced to a normal form through Fourier Integral Operator in [6]. In the same way, we will reduce it through a change of variable to have a model problem which can be understand as if the curve $\mathrm{E}$ was a line. Our objective is two-fold: we aim at describing the evolution of these dispersive terms and we want to consider initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
1.3 Main result
Instead of computing the asymptotics of the wave function itself, we focus on the probability densities such as the position one defined by $\left|\psi^{\varepsilon}_{t}(x)\right|_{\mathbb{C}^{2}}^{2}\mathrm{d}x$. Its physical interpretation is the probability of finding the particle at time $t$ in the position $x$. This approach generalizes the previous strategy and catches the information on the quantities that are not negligible in $\mathrm{L}^{2}$-norm. More precisely, we want to describe the evolution as $\varepsilon$ goes to 00 of the quadratic quantities
$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\coloneqq\int_{\mathbb{R}^{2}}\left\langle a(x)\psi^{\varepsilon}_{t}(x),\psi^{\varepsilon}_{t}(x)\right\rangle_{\mathbb{C}^{2}}\mathrm{d}x,$ (3)
where $\mathbb{C}^{2,2}$ is the set of $2$ by $2$ matrices and the scalar product of $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$ is $\left\langle v,w\right\rangle_{\mathbb{C}^{2}}\coloneqq v\cdot\overline{w}$.
For example, if $a$ is equal to a projector on a direction $\mathbb{R}\vec{V}$ for $\vec{V}\in\mathbb{R}^{2}\setminus\{0\}$ on a subspace $\Omega$, then $\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ describes the probability that $\psi^{\varepsilon}_{t}$ is collinear to $\vec{V}$ at time $t$ on the subspace $\Omega$.
We now introduce our assumptions on the function $\mathrm{m}$. Physically, the assumptions describe the interface between the two topological insulators.
Assumption 1.2.
@@ 1.3 Main result
To simplify the presentation and technical details, we will work with a global tubular neighborhood of the curve $\mathrm{E}$. To ensure the existence of such a neighborhood, we assume the following assumption. Actually, if we consider a neighborhood of a point of the curve $\mathrm{E}$, the local version of the following assumption will always be satisfied.
@@
Let $x_{0}\in\mathrm{E}$. Let us consider the parametrization $\gamma$ of the curve $\mathrm{E}$ and the vector-valued function $\nu$ such that $\gamma(0)=x_{0}$ and for all $s\in\mathbb{R}$,
@@
$\gamma^{\prime}(s)\coloneqq\frac{-{\nabla\mathrm{m}(\gamma(s))}^{\perp}}{\left|\nabla\mathrm{m}(\gamma(s))\right|}\;\;,\;\;\nu(s)\coloneqq\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|},$
where denotes the $\pi/2$-counterclockwise rotation. Then the curvature of the curve $\mathrm{E}$ at a point $\gamma(s)$ is the real number $\kappa(s)$ given by Frenet’s formula
$\kappa(s)\coloneqq\gamma^{\prime\prime}(s)\cdot\nu(s).$

### v3

# Propagation of Semiclassical Measures Between Two Topological Insulators

@@ Abstract
We study propagation in a system consisting of two topological insulators without a magnetic field, whose interface is a non-compact, smooth, and connected curve without boundary. The dynamics are governed by an adiabatic modulation of a Dirac operator with a smooth, effective variable mass. We determine the evolution of the semiclassical measure of the solution using a two-scale Wigner measure method, after reducing the Hamiltonian to a normal form.
Keywords: propagation, Dirac operator, semiclassical analysis, Wigner measures, semiclassical measures valued in a separable Hilbert space, time-dependent two-scale semiclassical measures.
@@ 1 Introduction
Topological insulators and Dirac-type equations.
Topological insulators are electronic materials that possess a bulk band gap, like ordinary insulators, but exhibit protected conducting states at their edges. The Dirac equation arises naturally in the context of topological insulators for honeycomb structures Fefferman et al. (2016b, a); Drouot (2019), such as graphene Raghu and Haldane (2008); Lee-Thorp et al. (2019); Drouot and Weinstein (2020).
@@
In this paper, we study the following system of Dirac evolution equation on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
@@
$\left\{\begin{array}[]{ll}\left(\varepsilon\mathrm{D}_{t}+\begin{pmatrix}\mathrm{m}(x)&\varepsilon\mathrm{D}_{1}-\mathrm{i}\varepsilon\mathrm{D}_{2}\\ \varepsilon\mathrm{D}_{1}+\mathrm{i}\varepsilon\mathrm{D}_{2}&-\mathrm{m}(x)\end{pmatrix}\right)\psi^{\varepsilon}(t,x)=0,&\forall(t,x)\in\mathbb{R}\times\mathbb{R}^{2},\\ \psi^{\varepsilon}(0,x)=\psi^{\varepsilon}_{0}(x),&\forall x\in\mathbb{R}^{2},\end{array}\right$. (1.1)
@@
where ${\mathrm{D}_{\#}\coloneqq-\mathrm{i}\partial_{\#}}$ for $\in\{t,1,2\}$ with $\partial_{j}\coloneqq\partial_{x_{j}}$ for $j\in\{1,2\}$, and where ${\varepsilon}$ is a small positive parameter. The family ${\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}}$ is uniformly bounded (with respect to $\varepsilon$) in ${\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$. The function $x\mapsto\mathrm{m}(x)$ is a smooth real-valued function, that is, $\mathrm{m}\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$, whose derivatives are all bounded. The interface between the two topological insulators is defined by
$\mathrm{E}\coloneqq\left\{x\in\mathbb{R}^{2}\mid\mathrm{m}(x)=0\right\}.$
To ensure that $\mathrm{m}$ separates $\mathbb{R}^{2}$ into two distinct regions, we assume that $\mathrm{E}$ is a smooth, connected curve without boundary. Under these conditions, there exists a unique family ${\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}}$ of solutions in ${\mathscr{C}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}$ to (1.1).
Physical interpretation.
The conducting states at the edge of topological insulators are called edge states. Their propagation is one of the most important problems in the physical applications of topological insulators and superconductors Bernevig (2013); Volovik (1989), photonics Lu et al. (2014); Raghu and Haldane (2008); Rechtsman et al. (2013), acoustics Peano et al. (2015); Ma et al. (2019), and fluid mechanics Graf et al. (2021). These states are physically realized through the combined effects of spin–-orbit interactions and time-reversal symmetry. More precisely, two-dimensional topological insulators are quantum spin Hall insulators, which are closely related to the integer quantum Hall state Hasan and Kane (2010). In such materials, there exists an energy gap between the valence and conduction bands; however, unlike in a trivial insulator, these bands are ”twisted” or ”crossing” Zhu et al. (2012). From a global point of view, Dirac-type equations often provide a simpler continuum macroscopic description of transport in a narrow energy band near a band crossing Bernevig (2013); Fruchart and Carpentier (2013); Volovik (1989). Such models typically describe the evolution of the electronic wave function in topological insulators, such as graphene monolayers subjected to an external field or to non-homogeneous deformations.
Here, the topological insulators under consideration are characterized by the sign of $\mathrm{m}$: the region where $\mathrm{m}$ is positive corresponds to a topological insulator of index $1$, while the region where $\mathrm{m}$ is negative corresponds to a topological insulator of index 1-1. The function $\psi^{\varepsilon}(t,\cdot)$, also denoted by $\psi^{\varepsilon}_{t}$, is interpreted as the wave function of an electron at time $t$ in a material consisting of two topological insulators.
Figure 1: Illustration of asymmetric transport for the mass $\mathrm{m}(x)=\arctan(x_{1})-x_{2}$.
Since equation (1.1) originates from quantum mechanics through the quantum spin Hall effect, we are naturally led to study the problem in phase space $\mathrm{T}^{*}\mathbb{R}^{2}=\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$. In this setting, $\xi\in\mathbb{R}^{2}$ is interpreted as the momentum of the particle.
@@ 1.1 Edge states generated by coherent states
The system (1.1) has already been studied in Bal et al. (2021), where the computed edge mode is generated by a very specific semiclassical wave packet. Following Drouot (2022), we define a wave packet concentrated at $(x_{0},\xi_{0})\in\mathbb{R}^{4}$, a point in phase space, as an $\varepsilon$-dependent function satisfying, for all $\varepsilon>0$,
@@
$\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}](x)\coloneqq\frac{e^{\frac{\mathrm{i}}{\varepsilon}\xi_{0}\cdot(x-x_{0})}}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),\quad x\in\mathbb{R}^{2},$
@@
with $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$. We say that $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$ is oriented along $\vec{V}\in\mathbb{C}^{2}$ if there exists $f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}\right)$ such that, for all $x\in\mathbb{R}^{2},\vec{f}(x)=f(x)\vec{V}$.
Geometrical setup.
We first introduce the geometrical setup.
Assumption 1.1 (Transversality condition).
@@
Let $x_{0}\in\mathrm{E}$. In view of Assumption 1.1, we consider an arc-length parametrization $\mathbf{t}$ of the curve $\mathrm{E}$, such that $\mathbf{t}(0)=x_{0}$, so
@@
$\mathbf{t}^{\prime}(s)\coloneqq\frac{{\nabla\mathrm{m}(\mathbf{t}(s))}^{\perp}}{\left|\nabla\mathrm{m}(\mathbf{t}(s))\right|},\quad s\in\mathbb{R},$
where denotes the counterclockwise rotation by $\pi/2$. We also introduce the vector-valued function $\mathbf{n}$ defined by